$SPAD/src/input f02bbf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
ia:=4
n:=4
alb:=-2.0
ub:=3.0
m:=3
iv:=4
a:Matrix SF:=
 [[0.5 ,0.0 ,2.3 ,-2.6 ],_
  [0.0 ,0.5 ,-1.4 ,-0.7 ],_
  [2.3 ,-1.4 ,0.5 ,0.0 ],_
  [-2.6 ,-0.7 ,0.0 ,0.5 ]]
result:=f02bbf(ia,n,alb,ub,m,iv,a,-1)

———-

References^{[1] nothing}

4